# Bidirectional Generative Pre-training for Improving Time Series Representation Learning

### v1

@@ 3.2 BiTimelyGPT Architecture
As shown in Figure 3.c, BiTimelyGPT employs the BAAR framework to alternately model left-to-right and right-to-left information flows across layers. In the odd-indexed layers $l\in\{1,3,\ldots,L-1\}$, BiTimelyGPT utilizes a forward retention to model left-to-right information with Equation 4, where the lower triangular decay matrix $\overrightarrow{{\bm{D}}}$ maintains this direction. In the even-indexed layers $l=(2,4,\ldots,L)$, BiTimelyGPT utilizes a backward retention to model right-to-left information:
@@
where a upper triangular decay matrix $\overleftarrow{{\bm{D}}}$ maintains right-to-left direction. Consequently, BiTimelyGPT alternates between forward and backward retention across layers, pre-training deep bidirectional representations.
Table 1: Model comparison from various perspectives. BiTimelyGPT achieves data preservation, bidirectionality, and full-rankedness.
Model Pre-training Data Preservation Bidirectionality Full-rankedness
PatchTST
Masking-based
CRT
Dropping-based
TimelyGPT
Next-Token Prediction
TimelyBERT
Masking-based
BiTimelyGPT Next-Previous-Token Prediction
In BiTimelyGPT, bidirectional generative pre-training aims to learn deep bidirectional representations from unlabeled data, preserving the original distribution and time-series shapelets without any data alterations. To maximize the joint probability of observing the entire sequence, BiTimelyGPT predicts the next token in the Layer $L-1$ and the previous token in the Layer $L$, respectively. We define the objective of Next-Previous-Token Prediction task as follows.
$\sum_{t=1}^{T}$ $\log p(x_{t}^{(L-1)}\mid x_{1}^{(L-1)},\ldots,x_{t-1}^{(L-1)})$
$+\log p(x_{t}^{(L)}\mid x_{t+1}^{(L)},\ldots,x_{T}^{(L)})$ (9)
The output projection layer takes each token’s representations $\overrightarrow{h}_{t}^{(L-1)}$ and $\overleftarrow{h}_{t}^{(L)}$ to predict the next and previous tokens, respectively. The pre-training loss is mean squared error (MSE) for continuous signals (e.g., biosignal) or cross-entropy for discrete signals (e.g., diagnosis codes). The [SOS] token in the final layer is used for fine-tuning as it aggregates a sequence representation for downstream tasks.
@@ 3.3 Full-rankedness in BiTimelyGPT
The self-attention mechanism often suffers from the low-rank bottleneck, leading to a loss of expressiveness in the rank-deficient attention matrix $\overleftrightarrow{{\bm{A}}}$ (Dong et al., 2021; Bhojanapalli et al., 2020). For single-head attention in Equation 1, when the dimension $d$ is less than the sequence length $N$, we have $\mathrm{rank}(\mathbf{Q}\mathbf{K}^{\top})\leq d$, resulting in a rank-deficient $\overleftrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ without Softmax activation. This bottleneck persists even with non-linear Softmax operation (Wang et al., 2020; Zhang et al., 2022b). This challenge is exacerbated in multi-head attention, where the dimension per head is much smaller than $N$ (Bhojanapalli et al., 2020). In contrast, GPT’s masked attention matrix $\overrightarrow{{\bm{A}}}$ in Equation 2 is full-rank (i.e., $\mathrm{det}(\overrightarrow{{\bm{A}}})\neq 0$), forming a lower triangular matrix with strictly positive diagonal elements post-Softmax activation. Although this full-rank matrix $\overrightarrow{{\bm{A}}}$ offers expressive representation power, it is limited to unidirectional contexts.
@@
BiTimelyGPT effectively overcomes the low-rank bottleneck by alternating between forward and backward retention, resulting in triangular retention matrices with non-zero diagonals. Specifically, the forward retention matrix $\overrightarrow{{\bm{R}}}$ in Equation 4 becomes full-rank when its diagonal elements are non-zero, ensuring $\mathrm{det}(\overrightarrow{{\bm{R}}})\neq 0$. For a diagonal element $\overrightarrow{{\bm{R}}}_{mm}={\bm{W}}_{Q}{\bm{X}}_{m}({\bm{W}}_{K}{\bm{X}}_{m})^{\top}$, this condition is satisfied by initializing these embeddings to prevent orthogonality and avoid all-zeros across the $d$ dimensions. This reasoning also applies to the backward retention $\overleftarrow{{\bm{R}}}$ in Equation 8. Consequently, both forward and backward retention matrices in BiTimelyGPT are full-rank, enabling to learn bidirectional contexts using more expressive attention matrices.
4 Experiment
4.1 Experiment Design
To evaluate BiTimelyGPT, we conducted both qualitative and quantitative analyses. Qualitatively, we visualized and interpreted three key advantages of BiTimelyGPT: data preservation during pre-training (Section 4.2), bidirectionality (Section 4.3), and full-rankedness (Section 4.4). For the quantitative analysis in Section 4.5, we assessed the performance of BiTimelyGPT on discriminative classification and regression tasks against a broad range of baseline models.
@@ 4.1 Experiment Design
Datasets. We used three publicly available, large-scale time-series datasets for pre-training: (1) the Sleep-EDF dataset with 7 types of biosignals across 1.2 billion timesteps (Kemp et al., 2000); (2) the PTB-XL dataset with 12 types of electrocardiogram data totaling 109 million timesteps (Alday et al., 2020); (3) the PPG-Dalia dataset with 4 types of photoplethysmograph data from 16.6 million timesteps (Reiss et al., 2019a). Additionally, we employed five other datasets for downstream discriminative tasks, namely Epilepsy (Andrzejak et al., 2001), EMG (Goldberger et al., 2000), RR, HR, SpO2 (Tan et al., 2021). The description and statistics of datasets are available in Section B.1.
@@
Qualitative Analysis Design. We conducted visualization analyses using the large-scale Sleep-EDF dataset, including (1) the gradient-based saliency for data preservation during generative pre-training; (2) the attention matrices for bidirectionality effectiveness; (3) the rank analysis of attention matrices for model expressiveness. To ensure consistency in comparison, we used the same patch-tokenizer with a non-overlapping patch size of 8, yielding $N=375$ tokens per 3000-timestep sequence (Nie et al., 2023).
Figure 4: Visualization of different pre-training methods. a. Distribution of average saliency scores within the extracted shaplets of sequences. b. A case example of time-series and its shapelet. c. Generative pre-training of BiTimelyGPT. The green rectangle indicates the shapelet discriminative region. d. Masking-based pre-training of PatchTST. e. Dropping-based pre-training of CRT.
Figure 5: Visualization of bidirectionality advantage for a discriminative shaplet. a. A case example of time-series and its data shapelet. b. The combined retention matrices from the last two layers of BiTimelyGPT. The green rectangle indicates the shapelet discriminative region. c. TimelyGPT’s causal retention matrix. d. GPT-2’s causal attention matrix. e. TimelyBERT’s bidirectional retention matrix.
@@
Quantitative Experiment Design. For classification, we pre-trained on the Sleep-EDF and PTB-XL datasets separately. Subsequently, the PTMs were fine-tuned on the Sleep-EDF, Epilepsy, PTB-XL, and EMG datasets. For regression, we pre-trained on both the PTB-XL and PPGDalia datasets, and then fine-tuned PTMs on the IEEEPPG, RR, HR, and SpO2 datasets. We divided each dataset into training (80%), validation (10%), and test (10%) sets. When the same dataset was used for both pre-training and fine-tuning, we fine-tuned the PTMs using 20% of the training set. All models adopted the preferred tokenization setup. We used accuracy and MAE as metrics for classification and regression, respectively.
Pre-training and Fine-tuning. BiTimelyGPT performed pre-training with Next-Previous-Token Prediction task, with fine-tuning on the [SOS] token in the final layer. PatchTST adopted a masking-based approach, masking 40% of its patches as zero (Nie et al., 2023). CRT utilized a dropping-based pre-training, discarding up to 70% of patches (Zhang et al., 2023). For transformers without established pre-training methods, we used a masking-based method by randomly masking 40% of timesteps (Zerveas et al., 2021). All transformers performed 20 epochs of pre-training with MSE loss, followed by 5 epochs of end-to-end fine-tuning.
@@
Bidirectional TimelyBERT Baseline. We also evaluated an intuitive baseline TimelyBERT, which uses bidirectional retention matrix $\overleftrightarrow{{\bm{R}}}$ in each layer. This baseline modifies the decay matrix as $\overleftrightarrow{{\bm{D}}}=\gamma^{|n-m|}$ in Equation 10. However, the TimelyBERT baseline faces two limitations: (1) this encoder-only model utilizes masking-based pre-training, resulting in distribution shift; (2) its bidirectional retention $\overleftrightarrow{{\bm{R}}}$ suffers from the low-rank bottleneck:
@@
4.2 Visualization of Different Pre-training Strategies
@@
We examined the impact of different pre-training methods on representation learning with respect to data preservation, focusing on BiTimelyGPT, masking-based PatchTST, and dropping-based CRT. We assessed the saliency of the $N\times d$ output embedding matrix generated during pre-training by computing the absolute value of the corresponding gradients with respect to the prediction target (Wagner & Foster, 2023). To identify discriminative patterns in time-series data (Ye & Keogh, 2009; Grabocka et al., 2014), we extracted shapelets from 10% of the Sleep-EDF dataset using the pyts package (Faouzi & Janati, 2020), followed by computing the average saliency values within these positions.
@@
As shown in Figure 4.a, BiTimelyGPT had higher saliency scores on the regions of data shapelets, suggesting its superior ability to uncover discriminative patterns compared to the baselines. Furthermore, we conducted a case study on a specific sequence with a distinct shapelet (Figure 4.b). The saliency heatmap of the output embedding matrix (Figure 4.c) revealed that BiTimelyGPT efficiently captured discriminative information surrounding the shapelet. In comparison, PatchTST (Figure 4.d) and CRT (Figure 4.e) displayed high saliency beyond the region of shapelet driven by the minimization of the reconstruction loss, failing to concentrate on discriminative patterns. These observations underscore that BiTimelyGPT’s pre-training effectively discerns discriminative information compared to masking-based and dropping-based approaches.

### v3

@@ 3.2 BiTimelyGPT architecture
As shown in Figure 3.c, BiTimelyGPT employs the BAAR framework to alternately model left-to-right and right-to-left information flows across layers. In the odd-indexed layers $l\in\{1,3,\ldots,L-1\}$, BiTimelyGPT utilizes the forward Retention to model left-to-right information with Equation 4, where the lower triangular decay matrix $\overrightarrow{{\bm{D}}}$ maintains this direction. In the even-indexed layers $l=(2,4,\ldots,L)$, BiTimelyGPT utilizes the backward Retention to model right-to-left information:
@@
where an upper triangular decay matrix $\overleftarrow{{\bm{D}}}$ maintains right-to-left direction. Consequently, BiTimelyGPT alternates between forward and backward Retention across layers, pre-training deep bidirectional representations.
In BiTimelyGPT, bidirectional generative pre-training aims to learn deep bidirectional representations from unlabeled data, preserving the original distribution and time-series shapelets without any data alterations. To maximize the joint probability of observing the entire sequence, BiTimelyGPT predicts the next token in Layer $L-1$ and the previous token in Layer $L$. We define the objective of the Next-Previous-Token Prediction task as follows.
$\sum_{t=1}^{T}$ $\log p(x_{t}^{(L-1)}\mid x_{1}^{(L-1)},\ldots,x_{t-1}^{(L-1)})+\log p(x_{t}^{(L)}\mid x_{t+1}^{(L)},\ldots,x_{T}^{(L)})$ (9)
The output projection layer takes each token’s representations $\overrightarrow{h}_{t}^{(L-1)}$ and $\overleftarrow{h}_{t}^{(L)}$ to predict the next and previous tokens, respectively. The pre-training loss is mean squared error (MSE) for continuous signals (e.g., biosignals) or cross-entropy for discrete signals (e.g., diagnosis codes). The [SOS] token in the final layer is used for fine-tuning as it aggregates a sequence representation for downstream tasks.
Table 1: Comparison of models based on BiTimelyGPT’s three key advantages: data preservation, bidirectionality, and full-rankedness.
Model Pre-training Data Preservation Bidirectionality Full-rank
PatchTST Masking-based
CRT Dropping-based
TimelyGPT Next-Token Pred
TimelyBERT Masking-based
BiTimelyGPT Next-Previous-Token Pred
@@ 3.3 Full-rankedness in BiTimelyGPT
The self-attention mechanism often suffers from the low-rank bottleneck, leading to a loss of expressiveness in the rank-deficient attention matrix $\overleftrightarrow{{\bm{A}}}\in\mathbb{R}^{N\times N}$ (Dong et al., 2021; Bhojanapalli et al., 2020). For the single-head attention in Equation 1, when the dimension $d$ is less than the sequence length $N$, we have $\mathrm{rank}(\mathbf{Q}\mathbf{K}^{\top})\leq d$, resulting in a rank-deficient $\overleftrightarrow{{\bm{A}}}$ without Softmax activation. This bottleneck persists even with non-linear Softmax operation (Wang et al., 2020; Zhang et al., 2022c). This challenge is exacerbated in multi-head attention, where the dimension per head is much smaller than $N$ (Bhojanapalli et al., 2020). In contrast, GPT’s masked attention matrix $\overrightarrow{{\bm{A}}}$ in Equation 2 is full-rank (i.e., $\mathrm{det}(\overrightarrow{{\bm{A}}})\neq 0$), forming a lower triangular attention matrix with strictly positive diagonal elements post-Softmax activation. Although this full-rank matrix $\overrightarrow{{\bm{A}}}$ offers expressive representation power, it is limited to unidirectional contexts.
@@
BiTimelyGPT effectively overcomes the low-rank bottleneck by alternating between forward and backward Retention, resulting in triangular Retention matrices with non-zero diagonals. Specifically, the forward Retention matrix $\overrightarrow{{\bm{R}}}$ in Equation 4 becomes full-rank when its diagonal elements are non-zero, ensuring $\mathrm{det}(\overrightarrow{{\bm{R}}})\neq 0$. For a diagonal element $\overrightarrow{{\bm{R}}}_{mm}={\bm{W}}_{Q}{\bm{X}}_{m}({\bm{W}}_{K}{\bm{X}}_{m})^{\top}$, this condition is satisfied by initializing these embeddings to prevent orthogonality and avoid all-zero values across the $d$ dimensions. This reasoning also applies to the backward Retention matrix $\overleftarrow{{\bm{R}}}$ in Equation 8. Consequently, both forward and backward Retention matrices in BiTimelyGPT are full-rank, enabling the learning of bidirectional contexts using more expressive attention matrices.
3.4 Computational complexity
BiTimelyGPT with its efficient Retention mechanism achieves $O(N)$ complexity, in contrast to BERT and GPT, which incur $O(N^{2})$ complexity (Katharopoulos et al., 2020). The self-attention mechanisms in Equation 1 and Equation 2 introduce a complexity of $O(N^{2}d)$. This quadratic computational bottleneck prevents standard Transformer models from modeling long sequences (i.e., $N>>d$). In contrast, BiTimelyGPT achieves linear complexity by leveraging both forward and backward Retention mechanisms. For forward Retention, $\text{Ret}(X_{n})=Q_{n}S_{n},S_{n}=K_{n}^{\top}V_{n}+\gamma S_{n-1}$, where both $Q_{n}S_{n}$ and $K_{n}^{\top}V_{n}$ have $O(d^{2})$ complexity. By recursively updating over $N$ timesteps, the total complexity becomes $O(Nd^{2})$. This is similarly applied in backward Retention. Therefore, BiTimelyGPT ensures $O(2Nd^{2})$ complexity, dramatically reducing computing resources when $N>>d$.
4 Datasets
4.1 Biosignal datasets
@@
We used three publicly available, large-scale biosignal time-series datasets for pre-training: (1) the Sleep-EDF dataset with 7 types of biosignals across 1.2 billion timesteps (Kemp et al., 2000); (2) the PTB-XL dataset with 12 types of electrocardiogram data totaling 109 million timesteps (Alday et al., 2020); (3) the PPG-Dalia dataset with 4 types of photoplethysmograph data from 16.6 million timesteps (Reiss et al., 2019). Additionally, we employed five other datasets for downstream discriminative tasks, including Epilepsy (Andrzejak et al., 2001), EMG (Goldberger et al., 2000), RR, HR, SpO2 (Tan et al., 2021). The description and statistics of datasets are available in Section B.1.
4.2 PopHR dataset
Population Health Record (PopHR) was established to monitor population health in Montreal, Quebec, Canada (Shaban-Nejad et al., 2016; Yuan et al., 2018). This healthcare administrative data includes clinical records of 1.3 million patients using International Classification of Diseases (ICD) codes. We extracted irregularly sampled time series from these clinical records. Specifically, we converted ICD-9 codes to phenotype codes (PheCodes) using the expert-defined PheWAS catalog (Denny et al., 2013, 2010). To identify relevant PheCodes for each target chronic disease, we calculated the prevalence ratio of each PheCode in the true patients compared to the general population. PheCodes with a prevalence ratio greater than 5 were included in our experiment. If no PheCode reached this threshold, we selected the top 5 PheCodes with the highest prevalence ratios. We excluded patients with fewer than 10 PheCodes from the dataset. This resulted in an irregularly sampled time series dataset comprising 47,000 patients, 2.2 million records, and 47 unique PheCodes.
5 Experiments
5.1 Qualitative analysis design on biosignals
@@
To evaluate three key advantages of BiTimelyGPT, we conducted qualitative analyses and interpreted visualization results on the large-scale Sleep-EDF dataset: (1) gradient-based saliency analysis for data preservation during pre-training (Section 6.1); (2) attention matrices for the effectiveness of bidirectionality (Section 6.2); (3) rank analysis of attention matrices for model expressiveness (Section 6.3). To ensure consistency in comparison, we used the same patch-tokenizer with a non-overlapping patch size of 8, yielding $N=375$ tokens per 3000-timestep sequence (Nie et al., 2023).
5.2 Quantitative experiment design on biosignals
@@
For the quantitative analysis in Section 6.4, we assessed BiTimelyGPT’s performance on downstream classification and regression tasks. For classification, we pre-trained on the Sleep-EDF and PTB-XL datasets separately. Subsequently, the PTMs were fine-tuned on the Sleep-EDF, Epilepsy, PTB-XL, and EMG datasets. For regression, we pre-trained on both the PTB-XL and PPGDalia datasets and then fine-tuned PTMs on the IEEEPPG, RR, HR, and SpO2 datasets. We divided each dataset into training (80%), validation (10%), and test (10%) sets. When the same dataset was used for both pre-training and fine-tuning, 20% of the training set was used for fine-tuning. All models adopted the preferred tokenization setup. We used accuracy and MAE as metrics for classification and regression, respectively.
5.3 Experiment design on PopHR’s irregularly sampled time series
The PopHR database offers rule-based labels for three chronic phenotypes: congestive heart failure (CHF), chronic obstructive pulmonary disease (COPD), and diabetes. For downstream multi-label classification task in Section 5.3, we used cross entropy and area under precision recall curve (AUPRC) to evaluate pre-training and fine-tuning, respectively.
5.4 Baselines
For the classification and regression tasks on biosignal datasets, we compared BiTimelyGPT against a broad range of baselines, including transformer PTMs, consistency-based PTMs, and recurrent models. Among transformer models, we included TimelyGPT, TimelyBERT, CRT, PatchTST, AutoFormer (Wu et al., 2021), FedFormer (Zhou et al., 2022), TST (Zerveas et al., 2021), and TimesNet (Wu et al., 2023). Consistency-based PTMs such as TS-TCC (Eldele et al., 2021), TS2Vec (Yue et al., 2022), and TF-C (Zhang et al., 2022b) were also evaluated. We also assessed recurrent models like LSTM, BiLSTM, and ELMo. Additionally, to explore the efficacy of our proposed BAAR framework, we applied it to other autoregressive models, namely GPT-2 and LSTM. For PopHR’s irregularly sampled time series, we compared BiTimelyGPT with efficient PTMs from Section 6.4, including PatchTST, TST, CRT, TS-TCC, and TF-C. We also assessed algorithms designed for irregularly sampled time series, including mTAND (Shukla and Marlin, 2021), GRU-D (Che et al., 2018), SeFT (Horn et al., 2019), and RAINDROP (Zhang et al., 2022a). We described the architectures and parameters for BiTimelyGPT and baselines in Section B.2.
@@ 5.4 Baselines
We also included a masking-based PTM TimelyBERT, which uses bidirectional Retention matrix $\overleftrightarrow{{\bm{R}}}$ in each layer. This baseline modifies the decay matrix to $\overleftrightarrow{{\bm{D}}}=\gamma^{|n-m|}$:
@@
However, the TimelyBERT baseline faces two limitations: (1) it relies on masking-based pre-training, leading to distribution shift; and (2) its bidirectional Retention $\overleftrightarrow{{\bm{R}}}$ suffers from the low-rank bottleneck.
5.5 Pre-training and fine-tuning
BiTimelyGPT performed pre-training with the Next-Previous-Token Prediction task, with fine-tuning on the [SOS] token in the final layer. For comparison, PatchTST adopted a masking-based method, masking 40% of its patches to zero (Nie et al., 2023). CRT used a dropping-based method, discarding up to 70% of patches (Zhang et al., 2023). For transformers without established pre-training methods, we used a masking-based method by randomly masking 40% of timesteps (Zerveas et al., 2021). All transformers performed 20 epochs of pre-training, followed by 5 epochs of end-to-end fine-tuning.
6 Results
6.1 Qualitative analysis of different pre-training strategies
@@
We examined the impact of different pre-training methods on representational learning with respect to data preservation, focusing on BiTimelyGPT, masking-based PatchTST, and dropping-based CRT. We assessed the saliency of the $N\times d$ output embedding matrix generated during pre-training by computing the absolute value of the corresponding gradients with respect to the prediction target (Wagner and Foster, 2023). To identify discriminative patterns in time-series data (Ye and Keogh, 2009; Grabocka et al., 2014), we extracted shapelets from 10% of the Sleep-EDF dataset using the pyts package and computed the average saliency values within these positions (Faouzi and Janati, 2020).
@@
As shown in Figure 4.a, BiTimelyGPT has higher saliency scores on the regions of data shapelets, suggesting its superior ability to uncover discriminative patterns compared to the baselines. Furthermore, we conducted a case study on a specific sequence with a distinct shapelet (Figure 4.b). The saliency heatmap of the output embedding matrix (Figure 4.c) reveals that BiTimelyGPT effectively captured discriminative information surrounding the shapelet. In comparison, PatchTST (Figure 4.d) and CRT (Figure 4.e) display high saliency beyond the region of the shapelet driven by the minimization of the reconstruction loss, failing to concentrate on discriminative patterns. These observations underscored that BiTimelyGPT’s pre-training effectively discerns discriminative information compared to masking-based and dropping-based approaches.
Figure 4: Visualization of different pre-training methods. a. Distribution of average saliency scores within the extracted shapelets of sequences. b. A case example of a time-series and its shapelet. c. Generative pre-training of BiTimelyGPT. The green rectangle indicates the shapelet discriminative region. d. Masking-based pre-training of PatchTST. e. Dropping-based pre-training of CRT.
6.2 Qualitative analysis of bidirectionality
We investigated the effectiveness of bidirectionality in attention matrices in a downstream classification task. We compared the attention matrices of fine-tuned models of BiTimelyGPT, TimelyGPT (Song et al., 2023), GPT-2 (Radford et al., 2019), and TimelyBERT. For BiTimelyGPT, we combined its lower and upper triangular Retention matrices, $\overrightarrow{{\bm{R}}}^{L-1}$ and $\overleftarrow{{\bm{R}}}^{L}$, to demonstrate their bidirectional representations. The distribution of attention scores within the shapelet regions reveals that BiTimelyGPT achieves the highest scores on the discriminative regions (Figure 5.a). Given a case sequence, BiTimelyGPT (Figure 5.c) exhibits localized and concentrated attention on the key discriminative region, efficiently integrating bidirectional contexts. In contrast, TimelyGPT (Figure 5.d) exhibits more limited focus on the discriminative segment. GPT-2 (Figure 5.e) has a broader but less precise range of attention around these key features. TimelyBERT (Figure 5.f), despite having decay mechanisms in both directions, struggles to concentrate attention on the discriminative region. These findings highlighted the effectiveness of BiTimelyGPT in identifying discriminative patterns by modeling bidirectional contextualized representations.
Figure 5: Visualization of bidirectionality advantage. a. Distribution of attention scores within the shapelet regions. b. A case example of time-series and its data shapelet. c. The combined Retention matrices from the last two layers of BiTimelyGPT. The green rectangle indicates the shapelet discriminative region. d. TimelyGPT’s causal Retention matrix. e. GPT-2’s causal attention matrix. f. TimelyBERT’s bidirectional Retention matrix.
Figure 6: Normalized cumulative singular values of the attention matrices. A more uniformly distributed curve of singular values indicates a higher rank attention matrix.
6.3 Qualitative analysis of attention matrix rankedness
We assessed the expressiveness of the attention matrices in various bidirectional baselines, including PatchTST, CRT, and TimelyBERT. We applied singular value decomposition to the attention matrices across different layers and heads, visualizing the normalized cumulative singular values (Figure 6). BiTimelyGPT exhibits a more uniform distribution of singular values, effectively overcoming the low-rank bottleneck. In contrast, the singular value curves for other bidirectional baselines skew toward the upper left corner, suggesting that the largest few singular values account for most information in the attention matrices. Notably, TimelyBERT has the most long-tailed spectrum distribution, due to its Retention matrix $\overleftrightarrow{{\bm{R}}}=\hat{\bm{Q}}\hat{\bm{K}}^{\top}\odot\overleftrightarrow{{\bm{D}}}$ lacking softmax’s non-linearity. As a result, the more even distribution of singular values in BiTimelyGPT indicates more expressive representation power, breaking the low-rank bottleneck seen in other bidirectional baselines with skewed distributions.